Mr.cowtongue wan te d to give his children 21 video games based im on how many chores they did around the house all year . Dan did 124 chores, nike did 327 and Tom did 252. How could Mr.cowtongue award the video games using each method


1
Expert's answer
2021-03-29T13:59:03-0400

Total number of chores = 124+ 327+252 = 703 chores

Dan = (124/703) × 21 = 3.7 . He will give Dan 4 games

Nike = (327/703) × 21 = 9.7 . He will give Nike 10 games

Tom = (252/703) ×21 = 7.5 . He will give Tom 7 games